OpenStudy (anonymous):

Given a polynomial f(x), if (x + 3) is a factor, what else must be true? f(3) = 0 f(-3) = 0 f(0) = 3 f(0) = -3

OpenStudy (anonymous):

@Directrix ?

OpenStudy (anonymous):

Would I be setting the equation to 0?

OpenStudy (jiteshmeghwal9):

By remainder theorem\[f(x)=q(x)(x+3)\]If u put x=-3\[f(-3)=q(-3)(-3+3)\]

Directrix (directrix):

>>Would I be setting the equation to 0? Yes If (x + 3) is a factor, then x = -3

Directrix (directrix):

If (x + 3) is a factor and x = -3, then then x = -3 is a zero of the function. It is also an x-intercept and has coordinates (-3,0). @Everly Which option do you think is correct?

OpenStudy (anonymous):

Thank you that makes sense and f(0) = -3 seems to make sense to me

Directrix (directrix):

>> f(0) = -3 means that if x = 0, then y = -3.

Directrix (directrix):

We want to get this point: (-3,0).

Directrix (directrix):

x = -3, y = 0

OpenStudy (anonymous):

Oh f(-3) = 0 ?

Directrix (directrix):

Yes.

OpenStudy (anonymous):

Great! and it makes sense to me too, thank you so much!

OpenStudy (jiteshmeghwal9):

u must know that if (x+a) is a factor of f(x) then x+a=0=>x=-a is a zero of the function. f(-a)=0

Directrix (directrix):

You are welcome.